# Equations involving the modular $j$-function and its derivatives

### v1

@@ 4. Solvability of Certain Equations Involving Periodic Functions
Let $F(z,w)=\sum_{k=0}^{n}z^{k}f_{k}(w)$ be a polynomial with coefficients in $\mathcal{P}(w)$, where $f_{n}\neq 0$, and let $s\in\{0,1\}$. Then there exist $10\leq x_{0}<$ 1, $y_{0}\geq 1$, $\delta>0$, $N,E_{0},E_{1}\geq 0$ such that
$|F(r+sw,w)|>\delta\sum_{k=0}^{n-1}|f_{k}(w)||r+sw|^{k}$
for all $w\in\operatorname{\mathbb{H}}$, $r\in\operatorname{\mathbb{R}}$ such that $y_{0}\leq\operatorname{Im}(w)$, $\operatorname{Im}(w)^{N}\leq|r|$, and one of the following holds:
@@
First, we assume that $\operatorname{Im}(w)$ is sufficiently large so that $w$ is in the domain of each $f_{k}$, and at least that $\operatorname{Im}(w)\geq 1$. We also assume that $0\leq\mathrm{Re}(w)\leq 2$. In particular, by Lemma 4.3, as soon as $\operatorname{Im}(w)$ sufficiently large, $f_{n}$ has no zeroes nor poles at $w$.
Write
$F(z,w)=f_{n}(w)\prod_{k=1}^{n}(z-\beta_{k}(w))$
where the $\beta_{k}$’s are the roots of $F(z,w)$ as a polynomial in $z$. Now consider $F(z+sw,w)$ as a polynomial in $z$ with coefficients in $\mathcal{P}(w)$ equipped with the valuation $\operatorname{ord}_{w=i\infty}$: by looking at its Newton polygon, we see that there are $(e_{\ell},d_{\ell})\in\operatorname{\mathbb{Q}}^{2}$ and $\alpha_{\ell}\in\operatorname{\mathbb{C}}^{\times}$ for $\ell=1,\dots,n$ such that for every $\varepsilon>0$, for each $k$ there is $\ell$ such that
$\left|\frac{sw-\beta_{k}(w)}{\alpha_{\ell}w^{-d_{\ell}}\exp(2\pi ie_{\ell}w)}-1\right|<\varepsilon$
for $\operatorname{Im}(w)$ sufficiently large.
We fix $\varepsilon$ sufficiently small so that for each $k$, there is a unique tuple $(\alpha_{\ell},e_{\ell},d_{\ell})$ satisfying the above inequality, provided $\operatorname{Im}(w)$ is sufficiently large (note that the tuples may be repeated, so the index $\ell$ may not be unique). Now fix an enumeration of the tuples, and rearrange the roots so for each $k$, $\beta_{k}$ satisfies the above inequality for $(\alpha_{k},e_{k},d_{k})$, that is for $\ell=k$. We will then think of each $\beta_{k}$ as a function of $w$ (not necessarily continuous) satisfying the above inequality, and so we will write for instance $\beta_{k}(w)\sim\alpha_{k}w^{-d_{k}}\exp(2\pi ie_{k}w)$ for $\operatorname{Im}(w)\to+\infty$.
We shall now bound $|r+sw-\beta_{k}(w)|$, distinguishing three cases depending on $(e_{k},d_{k})$.
  • If $(e_{k},d_{k})\geq(0,0)$, let $N_{k}=s$. Then for $\operatorname{Im}(w)$ large
    $|r+sw-\beta_{k}(w)|>\frac{|r|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{6}\quad\text{for }|r|\geq\operatorname{Im}(w)^{N_{k}}.$
  • If $e_{k}=0$ and $d_{k}<0$, then the argument of $sw-\beta_{k}(w)$ converges to $\theta_{k}=\arg(\alpha_{k})-\frac{d_{k}\pi}{2}\mod 2\pi$ for $\operatorname{Im}(w)\to+\infty$. If $\theta_{k}\notin\pi\operatorname{\mathbb{Z}}$, and if $s\neq 0$ we also have $d_{k}\leq-1$, then there is $\delta_{1}>0$ such that for $\operatorname{Im}(w)$ large
    $|r+sw-\beta_{k}(w)|>\delta_{1}(|r+sw|+|\beta_{k}(w)|).$
    If $\theta_{k}\in\pi\operatorname{\mathbb{Z}}$, let $N_{k}=-2d_{k}$, except if $s\neq 0$ and $d_{k}>-1$, in which case let $N_{k}=2$. Then there is $\delta_{2}>0$ such that for $\operatorname{Im}(w)$ large
    $|r+sw-\beta_{k}(w)|>2\delta_{2}|r|\geq\delta_{2}(|r+sw|+|\beta_{k}(w)|)\quad\text{for }|r|\geq\operatorname{Im}(w)^{N_{k}}.$
  • If $e_{k}<0$, then $sw-\beta_{k}(w)\sim-\beta_{k}(w)$ for $\operatorname{Im}(w)\to+\infty$. In particular we have
    $\log|sw-\beta_{k}(w)|$ $\sim\log|\beta_{k}(w)|\sim-2\pi e_{k}\operatorname{Im}(w),$
    $\arg(sw-\beta_{k}(w))$ $\sim\arg(\alpha_{k})+\pi+2\pi e_{k}\mathrm{Re}(w)\mod 2\pi.$
    Therefore, letting $E_{0,k}=-\frac{1}{4\pi e_{k}}$, $E_{1,k}=-\frac{1}{\pi e_{k}}$, for $\operatorname{Im}(w)$ large
    $\left|r+sw-\beta_{k}(w)\right|$ $>\frac{|r|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{4}$ for $\text{for }\operatorname{Im}(w)\leq E_{0,k}\log|r|,$
    $\left|r+sw-\beta_{k}(w)\right|$ $>\frac{|\beta_{k}(w)|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{4}$ for $\text{for }E_{1,k}\log|r|\leq\operatorname{Im}(w).$
@@
    Moreover, we can choose $x_{0}$ so that $\arg(\alpha_{k})+2\pi e_{k}x_{0}$ and $\arg(\alpha_{k})+2\pi e_{k}(x_{0}+1)$ are not in $\pi\operatorname{\mathbb{Z}}$. Note that we can do this simultaneously for all $k$ such that $e_{k}<0$. In particular, there is $\delta_{3}>0$ so that for $\operatorname{Im}(w)$ large
    $\left|r+sw-\beta_{k}(w)\right|>\delta_{3}(|r+sw|+|\beta_{k}(w)|)\quad\text{for}\mathrm{Re}(w)\in\{x_{0},x_{0}+1\}.$
Now let $E_{0}=\min_{k}E_{0,k}$, $E_{1}=\max_{k}E_{1,k}$, $N=\max_{k}N_{k}$, and fix $y_{0}\geq 1$ so that the above inequalities are true for $\operatorname{Im}(w)\geq y_{0}$. Then there is $\delta>0$ such that
@@
for any $w\in\operatorname{\mathbb{H}}$, $r\in\operatorname{\mathbb{R}}$ as in the conclusion. ∎
@@
Under the assumptions and notation of Lemma 4.5, let $(e,d)$ be the minimum order of $\frac{f_{k}}{f_{n}}$ at $i\infty$ for $k=0,\dots,n-1$ and suppose that $e<0$.
Then for all $r\in\operatorname{\mathbb{R}}$ sufficiently large the function $F(r+sw,w)$ has $-e$ zeroes, counted with multiplicity, within the region
@@
where $M=N$ if $N>0$ and $M=1$ otherwise. Moreover, for $w\in\partial\Xi_{r}$ we have
@@
For the sake of notation, let $F_{r}(w)\coloneqq F(r+sw,w)$. First, we apply Lemma 4.5 to $F_{r}(w)$ and find relevant constants $x_{0}$, $y_{0}$, $\delta$, $N$, $E_{0}$, $E_{1}$. If $N>0$, let $M=\lceil N\rceil$, otherwise let $M=1$. Let $r$ be large enough so that $|r|\geq 1$, $E_{0}\log|r|\geq y_{0}$, and $E_{1}\log|r|\leq|r|^{\frac{1}{M}}$, hence
@@
for $w\in\partial\Xi_{r}$. We also take $r$ sufficiently large so that each $f_{k}(w)$ does not have poles in $\Xi_{r}$, so in particular $F_{r}(w)$ is holomorphic on $\Xi_{r}$.
Note in particular that $F_{r}(w)$ is never zero on the boundary of $\Xi_{r}$. We shall compute the integral of the logarithmic derivative $\frac{F_{r}^{\prime}}{F_{r}}$ of $F_{r}$ along such boundary.
Vertical sides. First, we observe that
@@
Since $f_{k}\in\mathcal{P}(w)$, we have that $f_{k}(w+1)\sim f_{k}(w)$ as $\operatorname{Im}(w)\to+\infty$. Likewise, $r+sw+1\sim r+sw$ for $r+sw\to\infty$. Therefore, we can choose $r$ large enough so that the last factor on the right hand side has modulus less than $\frac{\delta}{2}$ for all $k$ and for any $w$ on the boundary of $\Xi$. We then have
$\left|\frac{F_{r}(w+1)}{F_{r}(w)}-1\right|=\left|\frac{F_{r}(w+1)-F_{r}(w)}{F_{r}(w)}\right|\leq\frac{\delta}{2|F_{r}(w)|}\left(\sum_{k=0}^{n}|f_{k}(w)||r+sw|^{k}\right)<\frac{1}{2}$
for $w\in\partial\Xi$. We may now choose a branch of $\log$ in the disc around $1$ of radius $\frac{1}{2}$ and estimate the integral along the vertical sides as
$\left|\int_{E_{0}\log|r|}^{|r|^{\frac{1}{M}}}\left(\frac{F_{r}^{\prime}(x_{0}+1+iy)}{F_{r}(x_{0}+1+iy)}-\frac{F_{r}^{\prime}(x_{0}+iy)}{F_{r}(x_{0}+iy)}\right)dy\right|=\left|\left.\log\left(\frac{F_{r}(w+1)}{F_{r}(w)}\right)\right|_{w=x_{0}+iE_{0}\log|r|}^{w=x_{0}+i|r|^{\frac{1}{M}}}\right|<2.$
Bottom side. By the previous estimate on $\frac{F_{r}(w+1)}{F_{r}(w)}$, for $r$ sufficiently large,
[…]
Top side. We now constraint $w$ to the region $\operatorname{Im}(w)=|r|^{\frac{1}{M}}$, $x_{0}\leq\mathrm{Re}(w)\leq x_{0}+1$. We have
@@
Recall that $r\sim\zeta w^{M}$ for some power $\zeta$ of $i$ depending on $M$ and the sign of $r$. For simplicity, fix the sign of $r$, and assume that $\zeta=1$, so that $r\sim w^{M}$. Then
@@
In particular, by Lemma 4.5, we find that $F(w^{M}+sw,w)-F_{r}(w)=o(F_{r}(w))$ for $r\to+\infty$. In other words, $F_{r}(w)\sim F(w^{M}+sw,w)$. The latter function is in $\mathcal{P}(w)$, thus it has some order $(e^{\prime},d^{\prime})$ at $i\infty$, and we can immediately conclude that $e^{\prime}=e$.
Therefore, we find that there are $\alpha\in\operatorname{\mathbb{C}}^{\times}$ and $d^{\prime}\in\operatorname{\mathbb{Q}}$ such that for $r$ large enough
$\left|\frac{F_{r}(w)}{\alpha w^{-\ell d^{\prime}}\exp(2\pi iew)}-1\right|<\frac{1}{4}.$
@@
[…]
Thus for sufficiently large $r$ we have
[…]
Conclusion. Summing up the contributions from all sides, with the appropriate orientations, we obtain
[…]
By the Argument Principle, the integral on the left hand side must be the difference between the number of zeroes and poles of $F_{r}(w)$ inside $\Xi$ (in particular an integer) multiplied by $2\pi i$. Since $F_{r}(w)$ is holomorphic on $\Xi_{r}$, it must have $-e$ zeroes in $\Xi_{r}$, counted with multiplicity. ∎
@@
This follows from combining Propositions 4.1 and 4.6, where we use $r=m$ a large integer and $s=1$. Note that if $\frac{f_{k}}{f_{n}}(z)\to\infty$ as $z\to i\infty$ then $\frac{f_{k}}{f_{n}}(z)$ must have exponential growth at $i\infty$ for it is periodic and so has a $q$-expansion. ∎

### v3

@@ 4. Solvability of Certain Equations Involving Periodic Functions
There is a positive $C\in\operatorname{\mathbb{R}}$ such that in the region
$U=\{w\in\operatorname{\mathbb{H}}:\operatorname{Im}(w)\geq C\log|w|\},$
there are holomorphic functions $\beta_{1},\dots,\beta_{n}:U\to\operatorname{\mathbb{C}}$ such that $F(\beta_{k}(w),w)=0$ for all $w\in U$, and if $F(\beta,w)=0$, then $\beta=\beta_{k}(w)$ for some $k$.
Moreover, there are $\alpha_{k}\in\operatorname{\mathbb{C}}^{\times}$, $e_{k},d_{k}\in\operatorname{\mathbb{Q}}$ such that $\beta_{k}(w)\sim\alpha_{k}w^{-d_{k}}q^{e_{k}}$ for $w\to\infty$ in $U$, where $q=\exp(2\pi iw)$ and $w^{-d_{k}}=\exp(-2\pi id_{k}\log(w))$ for some holomorphic branch of $\log(w)$ on $U$.
Proof.
It suffices to prove the conclusion for $F$ irreducible as a polynomial over $\mathcal{P}(w)$.
Let $F^{(z)}=\frac{\partial F}{\partial z}$. Then there are $G,H\in\mathcal{P}(w)[z]$ such that $GF+HF^{(z)}=1$. We take $C$ large enough that by Lemma˜4.3, the coefficients of $F$, $F^{(z)}$, $G$, and $H$ are holomorphic in the region $U=\{w\in\operatorname{\mathbb{H}}:\operatorname{Im}(w)\geq C\log|w|\}$. In particular, $F(z,w_{0})$ and $F^{(z)}(z,w_{0})$ have no common roots for any $w_{0}\in U$, and so by the implicit function theorem, and because $U$ is simply connected, there are $m$ holomorphic functions $\beta_{1},\ldots,\beta_{m}:U\to\operatorname{\mathbb{C}}$ such that $F(\beta_{t}(w),w)=0$ for every $t$ and $w\in U$, and moreover taking distinct values at all $w\in U$, thus if $F(\beta,w)=0$, then $\beta=\beta_{k}(w)$ for some $k$.
Now fix some $k$. For every $w\in U$, there is some $t$ such that $|\beta_{k}(w)^{t}f_{t}(w)|\geq|\beta_{k}(w)^{\ell}f_{\ell}(w)|$ for every $\ell$, and since $F(\beta_{k}(w),w)=0$, there is also $h\neq t$ such that $|\beta_{k}(w)^{h}f_{h}(w)|\geq\frac{1}{n}|\beta_{k}(w)^{t}f_{t}(w)|$. Let $U_{t,h}$ be the region where those inequalities hold. We have in particular
$1\geq|\beta_{k}(w)|\sqrt[h-t]{\left|\frac{f_{h}(w)}{f_{t}(w)}\right|}\geq\sqrt[h-t]{\frac{1}{n}}$
Let $(e_{k},d_{k})=\frac{\operatorname{ord}_{w=i\infty}(f_{t}/f_{h})}{h-t}$. By Lemma˜4.3 combined with the above inequalities, there is $N>1$ such that whenever $w$ is sufficiently large in $U_{t,h}$ we have $N\geq\frac{|\beta_{k}(w)|}{|w^{-d_{k}}q^{e_{k}}|}\geq\frac{1}{N}$ for some fixed determination of $w^{-d_{k}}$ on $U$. Choose $N$ so that it works simultaneously for any possible pair $t$, $h$. By continuity of $\beta_{k}$, the numbers $d_{k}$, $e_{k}$ do not depend on $t$, $h$.
Let $S$ be the set of indices $t$ such that $\operatorname{ord}_{w=i\infty}(f_{t})+(te_{k},td_{k})$ reaches a minimum value $(e,d)$. By construction, $S$ contains at least two elements. Write
$F(w^{-d_{k}}q^{e_{k}}z^{\prime},w)=q^{e}w^{d}(G_{0}(z^{\prime})+G_{1}(z^{\prime},w))$
where now $G_{0}$ is a non-trivial polynomial in $z^{\prime}$ with $|S|\geq 2$ terms and constant coefficients, and $G_{1}$ has coefficients that tend to 0 for $w\to\infty$ in $U_{t,h}$. Since $\frac{\beta_{k}(w)}{w^{-d_{k}}q^{e_{k}}}$ is bounded and continuous, it must converge to a non-zero root $\alpha_{k}$ of $G_{0}(z^{\prime})$, thus $\beta_{k}(w)\sim\alpha_{k}w^{-d_{k}}q^{e_{k}}$, as desired. ∎
We now provide an estimate on the size of $F_{r}(w)$ that we can use on the boundary of $\Xi_{r}$.
Lemma 4.6.
There exist $0\leq x_{0}<1$, $C,y_{0}\geq 4$, $D,E_{0},E_{1}\geq 0$ (with $E_{0}$ possibly $\infty$), $\delta>0$ such that
$|F_{r}(w)|=|F(r+sw,w)|>\delta\sum_{k=0}^{n-1}|f_{k}(w)||r+sw|^{k}$
for all $w\in\operatorname{\mathbb{H}}$, $r\in\operatorname{\mathbb{R}}$ such that $\operatorname{Im}(w)\geq y_{0}$, $|r|\geq\operatorname{Im}(w)^{D}$, and one of the following holds:
@@
We work in a region $U=\{w\in\operatorname{\mathbb{H}}:0\leq\mathrm{Re}(w)\leq 2\wedge\operatorname{Im}(w)\geq y_{0}\}$ for some $y_{0}$ sufficiently large as determined by this proof. We start by taking $y_{0}$ large enough that by Lemma˜4.3, $f_{n}$ has neither zeroes nor poles at $w$. We also require that $|r|\geq C$ for some $C$ sufficiently large, again as determined by this proof.
By Lemma˜4.5, provided $y_{0}$ is sufficiently large, there are holomorphic functions $\beta_{1},\ldots,\beta_{n}:U\to\operatorname{\mathbb{C}}$ parametrising the roots of $F(z,w)=0$ as functions of $w$, and for $w\to\infty$ in $U$ we have
$\beta_{k}\sim\alpha_{k}w^{-d_{k}}q^{e_{k}}$
for some $\alpha_{k}\in\operatorname{\mathbb{C}}^{\times}$ and $d_{k},e_{k}\in\operatorname{\mathbb{Q}}$.
Since we are assuming $0\leq\mathrm{Re}(w)\leq 2$, we also have $\operatorname{Im}(w)\leq|w|\leq\operatorname{Im}(w)+2$. We require that $C\geq 4$, $y_{0}\geq 4$, so that we have the following simple inequalities:
$|r+sw|\geq\max\{|r|-2,s\operatorname{Im}(w)\}\geq\max\left\{\frac{|r|}{2},s\frac{|w|}{2}\right\}\geq\frac{|r+sw|}{4}.$
It follows, for instance, that $|w|\sim\operatorname{Im}(w)\to+\infty$ for $w\to\infty$ in $U$. We shall omit the specification ‘in $U$’ in the rest of this proof.
We shall now bound $|r+sw-\beta_{k}(w)|$, distinguishing multiple cases depending on $(e_{k},d_{k})$.
  • If $(e_{k},d_{k})\geq(0,0)$, then $\beta_{k}(w)$ converges to a finite value $\gamma_{k}$ for $w\to\infty$. In this case, we require $C$ (if $s=0$) and $y_{0}$ to be large enough that
    $|r+sw|\geq 4|\gamma_{k}|\geq 2|\beta_{k}(w)|.$
    This ensures that
    $|r+sw-\beta_{k}(w)|\geq\frac{|r+sw|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{4}.$
  • Suppose that $e_{k}=0$ and $d_{k}<0$. For $|r|\geq\operatorname{Im}(w)^{2\max\{1,-d_{k}\}}$, we have both $|r|\geq 4|sw|$ and $|r|\geq 4|\beta_{k}(w)|$ for $w$ sufficiently large, and so for $y_{0}$ large we get
    $|r+sw-\beta_{k}(w)|\geq\frac{|r|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{8}.$
  • If $e_{k}<0$, then $sw-\beta_{k}(w)\sim-\beta_{k}(w)$ for $w\to\infty$. Since $\mathrm{Re}(w)$ is bounded, we get
    $\mathrm{Re}(\log(sw-\beta_{k}(w)))=\log|sw-\beta_{k}(w)|\sim\log|\beta_{k}(w)|\sim-2\pi e_{k}\operatorname{Im}(w).$
    We first give bounds when $|r|$ is roughly at least $|\beta_{k}|^{2}$, and when $|\beta_{k}|$ is roughly at least $|r|^{2}$. More precisely, for $y_{0}$ sufficiently large, we have
    $\left|r+sw-\beta_{k}(w)\right|$ $>\frac{|r|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{8}$ $\text{for }|r|\geq e^{-4\pi e_{k}\operatorname{Im}(w)},$
    $\left|r+sw-\beta_{k}(w)\right|$ $>\frac{|\beta_{k}(w)|}{2}\geq\frac{|r+sw|+|\beta_{k}(w)|}{8}$ $\text{for }|r|\leq e^{-\pi e_{k}\operatorname{Im}(w)}.$
    For $w\to\infty$ we also have
    $\operatorname{Im}(\log(sw-\beta_{k}(w)))\sim\arg(\alpha_{k})+\pi+2\pi e_{k}\mathrm{Re}(w)\mod 2\pi.$
@@
    Now choose $x_{0}$ so that $\arg(\alpha_{k})+2\pi e_{k}x_{0}$ and $\arg(\alpha_{k})+2\pi e_{k}(x_{0}+1)$ are not in $\pi\operatorname{\mathbb{Z}}$, and so are different from $\arg(r)\in\pi\operatorname{\mathbb{Z}}$. Note that we can do this simultaneously for all $k$ such that $e_{k}<0$. In particular, there is $\delta_{k}>0$ such that, after taking $y_{0}$ sufficiently large, we have
    $\left|r+sw-\beta_{k}(w)\right|>\delta_{k}(|r+sw|+|\beta_{k}(w)|)\quad\text{for }\mathrm{Re}(w)\in\{x_{0},x_{0}+1\}.$
Now, if $e_{k}=0$ for some $k$, let $D$ be the maximum between the values $2d_{k}$ and $2$ for such $k$, otherwise we can take $D=0$. If $e_{k}<0$ for some $k$, let $E_{0}$ be the maximum of $4\pi e_{k}$ and let $E_{1}$ be the minimum of $\pi e_{k}$ for such $k$, otherwise let $E_{0}=+\infty$ and $E_{1}=0$. Under the above choices, there is $\delta>0$ such that
@@
for any $w\in\operatorname{\mathbb{H}}$, $r\in\operatorname{\mathbb{R}}$ satisfying the requirements in the conclusion. ∎
We now recall the Argument Principle, which plays a key role in the proof of Proposition˜4.8.
Theorem 4.7 (Argument Principle, see e.g. [Lan99, Chapter VI, §1, Theorem 1.5]).
Let $f$ be a meromorphic function on a complex domain $\Omega$. Let $C$ be a simple closed curve (positively oriented) which is homologous to 0 in $\Omega$ and such that $f$ has no zeroes or poles on $C$. Let $Z$ and $P$ respectively denote the number of zeroes and poles (counted with multiplicity) of $f$ in the interior of $C$. Then
$2\pi i(Z-P)=\oint_{C}\frac{f^{\prime}(z)}{f(z)}dz=\oint_{f\circ C}\frac{dz}{z}.$
In the proof of the following proposition, we will integrate $\frac{F_{r}^{\prime}}{F_{r}}$ along the boundary of $\Xi_{r}$ and use the above estimates to find a positive lower bound, thus proving the existence of zeroes of $F_{r}$ within $\Xi_{r}$. For a more geometric description, integrating $\frac{F_{r}^{\prime}}{F_{r}}$ computes how many times the image $F_{r}(z)$ winds around 0 while $z$ moves along $\partial\Xi_{r}$. The bounds below will determine a rough picture of $F_{r}(\partial\Xi_{r})$, as in Figure˜3, and in turn determine the number of zeroes, counted with multiplicity.
@@
Let $(e,d)$ be the minimum order of $\frac{f_{k}}{f_{n}}$ at $i\infty$ for $k=0,\dots,n-1$ and suppose that $e<0$. We work under the notation of Lemma˜4.6.
Then for all $r\in\operatorname{\mathbb{R}}$ sufficiently large the function $F(r+sw,w)$ has e-e zeroes, counted with multiplicity, within the region (see Figure˜2)
@@
where $M=\lceil D\rceil$ if $D>0$ and $M=1$ otherwise. Moreover, for $w\in\partial\Xi_{r}$ we have
@@
Figure 2. The region $\Xi_{r}$ highlighted by the striped background.
Figure 3. Visual representation of the action of $F_{r}$ on the boundary of a typical rectangle $\Xi_{r}$ for $F$ of the form $Az^{2}+Bz+C{j(z)}^{2}+Dj(z)+E$. The term $j^{2}$ has lowest order (2,0)(-2,0), so $F_{r}$ winds around 0 twice while following the top side of the rectangle.
@@
Recall that $F_{r}(w)=F(r+sw,w)$. First, we apply Lemma˜4.6 to $F_{r}(w)$ and find relevant constants $x_{0}$, $y_{0}$, $\delta$, $C$, $D$, $E_{0}$, $E_{1}$. Let $M=\lceil D\rceil$ if $D>0$ and $M=1$ otherwise. Let $r$ be large enough so that $|r|\geq C$, $E_{0}\log|r|\geq y_{0}$, and $E_{1}\log|r|\leq|r|^{\frac{1}{M}}$, hence
@@
for $w\in\partial\Xi_{r}$. It follows, for instance, that $F_{r}(w)$ has order $(e,d)$ at $i\infty$. We also take $r$ sufficiently large so that each $f_{k}(w)$ does not have poles in $\Xi_{r}$, so in particular $F_{r}(w)$ is holomorphic on $\Xi_{r}$.
Note that $F_{r}(w)$ is never zero on the boundary of $\Xi_{r}$, thus its logarithmic derivative $\frac{F_{r}^{\prime}}{F_{r}}$ is holomorphic there. We shall now compute the integral of $\frac{F_{r}^{\prime}}{F_{r}}$ of $F_{r}$ along such a boundary.
Vertical sides. We show that the images of the vertical sides of $\Xi_{r}$ under $F_{r}$ must be close to each other, and so their contributions cancel out as they are taken with opposite orientations. For instance, if $F_{r}$ happens to be $1$-periodic (as a function of $w$; for instance, when $s=0$ and the coefficients of $F$ are $1$-periodic) then the images of these vertical sides are identical.
First, we observe that
@@
Since $f_{k}\in\mathcal{P}(w)$, we have that $f_{k}(w+1)\sim f_{k}(w)$ as $\operatorname{Im}(w)\to+\infty$. Likewise, $r+sw+1\sim r+sw$ for $r+sw\to\infty$. Therefore, we can choose $r$ large enough so that the last factor on the right-hand side has modulus less than $\frac{\delta}{2}$ for all $k$ and for any $w$ on the boundary of $\Xi$. We then have
(4.1) $\left|\frac{F_{r}(w+1)}{F_{r}(w)}-1\right|=\left|\frac{F_{r}(w+1)-F_{r}(w)}{F_{r}(w)}\right|\leq\frac{\delta}{2|F_{r}(w)|}\left(\sum_{k=0}^{n}|f_{k}(w)||r+sw|^{k}\right)<\frac{1}{2}$
for $w\in\partial\Xi$. We may now choose a branch of $\log$ in the disc around $1$ of radius $\frac{1}{2}$ and estimate the integral along the vertical sides as [^4] Recall that $\frac{(fg)^{\prime}}{fg}=\frac{f^{\prime}}{f}+\frac{g^{\prime}}{g}$, and that $\int_{\gamma}\frac{f^{\prime}}{f}=\log(f(\gamma(1)))-\log(f(\gamma(0)))$ whenever $\gamma:[0,1]\to\operatorname{\mathbb{C}}$ is a path and $f\circ\gamma$ takes values in a simply connected region on which we have fixed a continuous branch of $\log$.
$\left|\int_{E_{0}\log|r|}^{|r|^{\frac{1}{M}}}\left(\frac{F_{r}^{\prime}(x_{0}+1+iy)}{F_{r}(x_{0}+1+iy)}-\frac{F_{r}^{\prime}(x_{0}+iy)}{F_{r}(x_{0}+iy)}\right)dy\right|=\\ =\left|\left.\log\left(\frac{F_{r}(w+1)}{F_{r}(w)}\right)\right|_{w=x_{0}+iE_{0}\log|r|}^{w=x_{0}+i|r|^{\frac{1}{M}}}\right|<\log\left(\frac{3}{2}\right)-\log\left(\frac{1}{2}\right)=\log(3)<2.$
Bottom side. We now show that the image of the bottom side is away from 0 and cannot wind much. Using (4.1), for $r$ sufficiently large,
$\left|\int_{x_{0}}^{x_{0}+1}\frac{F_{r}^{\prime}(x+iE_{0}\log|r|)}{F_{r}(x+iE_{0}\log|r|)}dx\right|=\left|\log\left(\frac{F_{r}(x+iE_{0}\log|r|+1)}{F_{r}(x+iE_{0}\log|r|)}\right)\right|<\log\left(\frac{3}{2}\right)<1.$
Top side. We show that $F_{r}$ behaves like $\exp(2\pi iew)$ on the top side of $\Xi_{r}$, and so the image under $F_{r}$ is roughly a circle traversed approximately $e$ times.
We now constraint $w$ to the region $\operatorname{Im}(w)=|r|^{\frac{1}{M}}$, $x_{0}\leq\mathrm{Re}(w)\leq x_{0}+1$. We have
@@
By construction, $r\sim\zeta w^{M}$ for some power $\zeta$ of $i$ depending on $M$ and the sign of $r$. For simplicity, fix the sign of $r$, and assume that $\zeta=1$, so that $r\sim w^{M}$. Then
@@
In particular, by Lemma˜4.6, we find that $F(w^{M}+sw,w)-F_{r}(w)=o(F_{r}(w))$ for $r\to+\infty$. Since $F(w^{M}+sw,w)$ is in $\mathcal{P}(w)$, it has an order $(e^{\prime},d^{\prime})$ at $i\infty$, and in fact $e^{\prime}=e$ because the term $w^{M}$ cannot alter the exponential growth.
Therefore, we find that there is $\alpha\in\operatorname{\mathbb{C}}^{\times}$ such that for $r$ large enough
$\left|\frac{F_{r}(w)}{\alpha w^{-d^{\prime}}\exp(2\pi iew)}-1\right|<\frac{1}{4}.$
@@
$\left|\int_{x_{0}}^{x_{0}+1}\frac{((x+i|r|^{\frac{1}{M}})^{-d^{\prime}}\exp(2\pi ie(x+i|r|^{\frac{1}{M}})))^{\prime}}{(x+i|r|^{\frac{1}{M}})^{-d^{\prime}}\exp(2\pi ie(x+i|r|^{\frac{1}{M}}))}dx-2\pi ie\right|=\left|\int_{x_{0}}^{x_{0}+1}\frac{d^{\prime}}{x+i|r|^{\frac{1}{M}}}dx\right|<1.$
Thus, for sufficiently large $r$ we have
$\left|\int_{x_{0}}^{x_{0}+1}\frac{F_{r}^{\prime}(x+i|r|^{\frac{1}{N}})}{F_{r}(x+i|r|^{\frac{1}{N}})}dx-2\pi ie\right|\leq\left|\left.\log\left(\frac{F_{r}(w)}{\alpha w^{-d^{\prime}}\exp(2\pi iew)}\right)\right|_{w=x_{0}+i|r|^{\frac{1}{N}}}^{w=x_{0}+1+i|r|^{\frac{1}{N}}}\right|+1<2.$
Conclusion. Using the above estimates we can now find the winding number of $F_{r}(\partial\Xi_{r})$ at 0.
Summing up the contributions from all sides, with the appropriate orientations, we obtain
$\left|\frac{1}{2\pi i}\oint_{\partial\Xi_{r}}\frac{F_{r}(w)^{\prime}}{F_{r}(w)}dw+e\right|<\frac{1}{2\pi}\left(2+1+2\right)<1.$
By the Argument Principle, the integral on the left-hand side must be the difference between the number of zeroes and poles of $F_{r}(w)$ inside $\Xi$ (in particular an integer) multiplied by $2\pi i$. Since $F_{r}(w)$ is holomorphic on $\Xi_{r}$, it must have e-e zeroes in $\Xi_{r}$, counted with multiplicity. ∎
@@
This follows from combining Propositions 4.1 and 4.8, where we use $r=m$ a large integer and $s=1$. Note that if $\frac{f_{k}}{f_{n}}(z)\to\infty$ as $z\to i\infty$ then $\frac{f_{k}}{f_{n}}(z)$ must have exponential growth at $i\infty$ for it is periodic and so has a $q$-expansion. ∎